OpenStudy (anonymous):

The equation of line XY is (y-3) =-2/3 What is the slope of a line perpendicular to XY? Select one: a. -2 b. 3 c.-2/3 d.3/2

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

you're missing the "x" somewhere, aren't you? Check your question again :)

OpenStudy (anonymous):

The equation of line XY is (y-3) = -2/3 (x - 4). What is the slope of a line perpendicular to XY? Select one: a. -2 b. 3 c.-2/3 d.3/2

OpenStudy (anonymous):

here

OpenStudy (anonymous):

Its currently in point-slope form, that's okay Point-Slope Equation of a Line y – y1 = m(x – x1), where m is the slope and (x1, y1) is a point on the line.

OpenStudy (kc_kennylau):

So now you have the slope, remember the relation between a line and its perpendicular line? :)

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

sorry, I have to go now, @Phaen will continue to help you if he wants to :)

OpenStudy (anonymous):

watever

OpenStudy (anonymous):

do you know the term negative reciprocal?

OpenStudy (texaschic101):

y - 3 = -2/3(x - 4) slope here is -2/3 perpendicular slope will have negative reciprocal slope. All that means is " flip " the slope and change the sign. So the negative reciprocal slope is 3/2 (see how I flipped the slope and changed the sign)